A Complete Bibliography of Publications in Synthese, 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

M′′ [15]. S5 [15].

A. [83]. Abstracts [209, 220, 263, 250, 273]. acquisition [243]. Actance [90]. action [280]. addition [213]. address [127]. adopted [41]. affects [91]. algebra [16]. Algebraic [106]. algèbre [16]. Allgemeinen [119]. Allocution [128]. Analogies [114]. Analyse [45]. analysis [80, 59, 180, 61, 45]. Analyticity [98, 240]. anwendungen [40]. any [253, 254]. applicable [194]. applications [40]. applied [17, 37, 66]. approach [151, 111]. Arbini [244]. Argument [135, 158]. Aristotle [105]. art [121]. article [143, 152]. artificial [92]. aspects [21, 81]. assert [176]. Association [28, 94]. assumptions [226, 286]. atomism [203]. Austin [166]. Axiomatic [212, 8]. Axiomatizability [20]. axioms [215].

B [79]. background [247]. background-knowledge [247]. Bare [6]. Bayesian [279]. be [90]. Beginning [225]. Behavioral [227]. Beiträge [45].

Beleuchtung [40]. belief [233]. Bemerkungen [172]. ’Beth [130, 130, 129, 138]. between [96, 269, 24]. biased [279]. Bibliography [138]. biological [267, 113, 202]. biology [114, 269, 264, 270, 271, 265, 266, 272]. body [63]. Book [87, 51, 62, 74, 57]. Bookreviews [10]. Books [140, 154, 182, 188, 199, 211, 221, 238, 251, 274, 282, 288]. Briefe [207]. British [263]. Bruno [254]. Bunge [245].

calculus [84, 107, 179]. calculuses [33]. can [90]. candor [115]. Carnap [40, 34]. cases [125]. Causal [193]. Causality [195, 117, 194]. Causation [285]. certain [131]. certaines [131]. Certains [21]. Changing [247]. chapter [224]. characterizability [175]. Choice [216, 145, 21, 148]. choix [21]. Church [52]. circuits [106]. class [118]. classes [131]. classification [73]. classique [135]. closely [41]. cognitive [286, 43]. Color [170]. Comment [162, 56]. Comments [243, 83, 174, 76, 258, 82, 75]. Communication [104, 91]. Communications [95, 144, 198]. comparison [26]. competence [243]. Complementarity [59]. complex [100]. Complexity [102]. computability [38]. concept [160, 53, 204, 259]. concepts [239, 77, 73, 89, 170]. Conceptual [161, 256, 162]. concerning [33]. Conclusions [136]. Confirmation [256, 192]. Congresses [9]. connection [193]. consistency [236]. Constructive [8]. contemporary [71]. contrasts [116]. contributions [156, 45]. controversies [280]. controversy [69]. convention [50]. Conventionalism [228]. corpuscles [24]. corpuscules [24]. Correct [246]. Correction [189]. Corrections [246, 245]. cost [150]. counterpart [37, 66]. covering [276]. Creative [107]. creativity [270]. criteria [227]. critics [206]. cultural [80, 55].

dans [50, 131]. datum [166]. decision [287, 146, 184]. deducing [90]. Deduction [3]. deductive [61]. defense [261, 271]. definability [43]. definition [204]. definitions [107]. demand [88]. demand-sentences [88]. description [113, 277, 217, 27]. Design [112]. determinism [89]. Deutscher [207]. developed [90]. development [55]. devious [101]. Did [205]. different [15, 53, 270]. différentes [15]. dimensions [53]. d’Interaction [24]. discontinuity [164, 163]. discovery [196]. Discussion [79, 254, 31, 254]. dispositions [170]. dispute [269]. division [95, 144, 198]. Dr. [76, 258, 75]. Duhem [71]. d’un [50]. d’une [21]. Dutch [28, 94].

E. [138]. econometric [284]. econometrics [285]. economic [285]. economics [283]. Editor [30]. Editorial [1, 126, 137, 155, 248, 252, 275, 222, 200]. Einige [45]. Einstein [206]. elementarity [159]. elementary [6]. empirical [226, 26]. Encyclopedia [44]. enduring [269]. Energy [213]. Entitled [176]. entre [24]. Entropy [184, 213, 60]. epistemological [58, 158]. Epistemology [32]. erkenntnistheoretischer [58]. Ernst [207, 202, 201, 207, 205]. Errata [29]. espace [50]. espaces [132]. estimation [4]. evaluation [71]. Evidence [2]. evolutionary [194]. Exact [28]. Exacte [28, 94]. existence [192]. experimental [12, 265]. Explanation [277, 280, 85, 278]. explanatory [266, 157]. Exposition [145]. expressions [41]. Extension [35, 131]. extensionality [115]. extensions [105]. extensive [212].

fallacy [46]. falsifiability [71]. few [130]. finding [48]. Finetti [254]. finies [131]. finite [131]. Fitness [266]. fonction [21]. fonctionnelle [24]. Formal [25, 80, 11, 97, 147]. formalised [96]. Formalized [92]. formation [77]. formulation [151]. Foundations [246, 236, 245]. Frazer [172]. Frazers [172]. French [50, 14, 134, 15, 16, 131, 254, 21, 130, 132, 133, 135, 24]. function [21]. functional [24]. functions [287]. Fundamental [215]. fuzziness [98].

game [111, 173, 174]. game-theoretical [111]. general [260, 38, 159, 119]. generalization [194]. genereller [260]. genetic [46]. geometric [133]. géométrique [133]. geometry [20, 169]. German [58, 40, 18, 260, 119, 45, 207, 259, 172]. Gibson [241]. Gödel [52]. good [43]. grading [147]. group [287]. Grünbaum [83]. Gültigkeit [260].

Hegel [69]. Henri [133]. hidden [81]. hidden-variable [81]. Hinsicht [58]. history [95, 144, 198, 44]. hypothèse [131]. hypotheses [276, 192]. hypothesi [157]. hypothesis [131].

iconic [18]. ideas [156]. ikonische [18]. Illumination [40]. impermissible [42]. implication [3]. inconsistencies [39]. incorrect [246]. indeterminacy [228]. Indian [25]. Individuality [270]. induction [120, 253, 258, 143, 257, 49, 254]. Inductive [125, 39, 255, 61]. inertia [67]. inexact [239]. inference [183, 61]. influences [55]. informantion [184]. Information [183, 60, 277, 54, 185]. Initial [253, 254]. innate [156]. innateness [157]. innovation [256]. insight [22]. insights [230]. intension [35]. intensional [64]. intentions [99]. interacting [5]. Interaction [24]. International [44, 95, 144, 198]. Interpretations [70]. Intrinsic [149]. introduction [222, 224]. introductions [152]. intuitionistic [132]. intuitionniste [132]. invariance [256]. invariant [255]. isolated [50]. isolé [50]. issues [280].

Journal [263].

Kierkegaard [69]. kinds [15]. knowledge [247, 202, 54, 152, 27].

language [243, 58, 96, 54, 77, 173, 179, 174]. language-game [173, 174]. languages [96, 33, 64, 92]. law [160, 67]. lawlikeness [258]. laws [90, 276]. learning [151, 109]. Letters [207]. levels [53]. linéaires [132]. linear [132]. linguistic [175, 243, 256]. Linguistically [255]. linguistics [157]. locutions [42]. Logic [28, 94, 235, 135, 179, 95, 144, 198, 32, 239, 255, 148, 33, 99, 88, 146, 25, 63, 134]. Logica [28, 94]. Logical [257, 59, 84, 190, 16, 21, 115, 34, 78, 49]. logics [15]. Logique [134, 16, 21, 135]. logiques [15]. logistics [40]. logistik [40].

Mach [203, 202, 201, 204, 207, 205]. macrostates [186]. makes [194]. Marcus [79, 65]. mass [204]. Material [124]. mathematical [73, 257]. mathematics [110, 8, 36, 37, 66, 26]. maximum [115]. meaning [178, 224, 56, 26]. meanings [41]. measurement [141, 212, 142, 159]. mécanique [135]. Mechanics [135, 59]. Mechanism [267, 269, 52]. Memorial [127]. Mergers [283]. metaphors [110]. Metaphysical [113]. method [265, 129]. methodology [267, 95, 144, 198]. methods [180]. Mew [236]. microphysics [14]. microphysique [14]. mind [63]. mind-body [63]. minds [122]. miss [205]. modal [84, 15]. modales [15]. Modalities [64]. modality [232]. Model [22, 27, 14, 19, 21, 48, 63, 24]. Modèle [24, 14, 21]. Modèles [15, 16]. Modelle [18]. Models [90, 17, 110, 11, 12, 151, 109, 23, 157, 56, 26, 147, 15, 16, 18]. Modification [236]. modular [63]. Molecular [269]. Molecules [23]. momental [90]. monism [167]. Mood [173, 174]. morphisms [213].

names [6]. natural [96, 112]. nature [141, 85, 142]. naturwissenschaftlichen [259]. Nederlandse [28, 94]. Neural [186]. Newton [206]. No [196]. Non [61, 43, 11, 107, 135]. Non-cognitive [43]. non-creative [107]. Non-deductive [61]. non-formal [11]. note [30, 137, 155, 248, 252, 275]. Notes [72, 150]. notion [14, 21]. numerical [4].

Object [224]. objects [80, 124, 242]. observateur [50]. observer [50]. Occam [37, 66]. ontological [37]. Operational [81]. operationism [214]. order [213]. ordinary [225]. organicism [269]. Organisation [45]. organization [102, 111, 45]. organizes [5]. other [122]. our [247].

paper [79, 83, 254, 76, 258, 75]. papers [82]. partial [71]. particular [131]. particulars [6]. particulières [131]. patterns [266]. perception [180, 201, 169]. perfect [185]. Performatives [177]. Permissible [42]. Philosophen [207]. Philosopher [207]. philosophical [31]. philosophie [133]. philosophies [116]. Philosophy [28, 94, 263, 95, 144, 198, 201, 264, 36, 133, 223, 283, 250, 273]. Physical [94, 134, 167, 13, 272]. Physics [246, 201, 245, 196, 19, 78]. physique [134]. Platonism [235]. Poincaré [133]. points [20, 164, 163]. Polanyi [268]. population [5]. pour [15, 135]. pragmatic [175]. predication [229]. prediction [277]. predictive [90]. preferability [149]. Preference [148, 216, 215, 150]. prerequisite [31, 253, 254]. prescribing [217]. present [36]. presuppositions [113]. principal [261]. principle [194, 49].

Principles [119, 147]. Prinzipien [119]. probabiliste [135]. Probabilistic [135]. probabilities [253, 190, 254]. Probability [109, 17, 143, 107, 257, 259]. Problem [45, 149, 122, 240, 4, 257, 259]. Problematik [259]. problems [61, 169, 233]. procedures [279]. processes [109]. Prof. [82]. Professor [65]. Program [204]. propositional [84]. propositions [6]. psychic [53]. psychology [12, 7]. pure [37, 66]. purpose [141].

qualities [242]. quality [286]. Quantifying [234]. quantique [135]. Quantum [135, 59, 117, 70, 81]. Quelques [130]. questionnaires [41]. Quine [226, 232, 230, 71, 236, 223].

rôle [50]. radical [227]. random [5]. rational [31]. razor [37, 66]. real [213]. realism [168, 241, 121]. reasoning [61]. reasons [168]. received [140, 154, 182, 188, 199, 211, 221, 238, 251, 274, 282, 288]. Rectification [210]. reduction [268, 272]. redundant [106]. reference [124]. refuted [166]. regions [192]. rejection [121]. related [41, 89]. relation [63]. relations [131, 216, 215]. relationship [96, 77]. Relativisation [21]. Relativitätstheorie [119]. Relativity [119, 195, 159]. Remark [172]. Remarks [69, 120, 130, 33, 132]. Remarques [132, 130]. Replies [237]. Reply [244, 65, 218]. requirements [195]. Review [103, 208, 93, 219, 281, 143, 152]. Reviews [187, 197, 165, 249, 139, 262, 181, 153, 171, 87, 51, 62, 74, 57]. revolution [85]. revolutions [162, 161]. roads [101]. role [50, 58]. Rolle [58]. Rudolf [40, 40]. rules [61]. Russell [118]. Ruth [79].

S [76, 82, 75]. sa [131]. Sätze [260]. saying [231]. Scaling [142]. Schiller [82]. Science [44, 121, 95, 144, 198, 201, 55, 116, 101, 272, 123, 223, 161, 48, 283, 263, 250, 273, 162]. sciences [11, 26, 28, 94]. scientific [85, 259]. scientist [7]. selection [112]. Semantic [175, 160, 66]. Semantics [13, 97]. sense [166]. sense-datum [166]. sensory [242]. sentences [260, 177, 88]. sequential [106]. Siegel [82]. signs [124]. simple [100]. simplicity [125]. simplification [106]. single [91]. Singular [229]. situation [36]. slight [236]. Social [55]. sociology [90]. Some [120, 33, 163, 169, 233, 147, 266, 21, 45]. sortes [15]. sources [160]. space [50]. spaces [132]. special [125]. specification [285]. Speculations [108]. speech [180]. Sprache [58]. statement [271]. statistical [276]. Statistics [258, 279, 257]. status [49]. Stenius [174]. Stephen [162]. stipulation [84]. stochastic [85]. Strategy [146]. strict [3]. Structural [123]. structure [268, 190, 78]. structures [216, 25]. study [11, 116, 41, 47]. subjective [259]. subjektiven [259]. supererogation [149]. survey [97]. syllogistic [105]. Symbolic [18]. Symbolische [18]. synonymy [72, 43]. syntactic [21]. syntactical [230]. syntactiques [21]. Systemen [45]. systems [113, 194, 45]. tableau [129]. tableau-method [129]. Tableaux [130]. Tables [130]. target [205]. targeted [45]. Teleologie [45]. teleology [45]. temperature [214]. terms [229, 63]. testable [90]. testing [151]. their [177]. theorem [261, 52]. Theoretical [271, 111]. théorie [134, 132, 24]. theories [108, 13, 286, 71, 100, 81, 206]. Theory [134, 267, 145, 166, 58, 202, 117, 156, 285, 148, 277, 195, 111, 132, 38, 152, 118, 4, 142, 159, 70, 48, 24, 119]. there [242]. thermodynamics [212]. thesis [71]. things [225]. thought [109]. three [53]. tools [284]. Total [71]. Toulmin [162]. Transacts [90]. transformations [86]. translation [224, 227, 228]. Truth [178, 84, 191, 34]. Two [116, 152, 224]. Typology [41].

uncertainty [145]. Unconventional [135]. Unified [44]. uniform [54]. Union [144, 95, 198]. universe [89]. unjust [214]. usage [175]. use [12, 177, 259]. uses [37, 56, 26, 66]. utility [150].

valid [253, 254]. validity [260]. value [185]. variable [81]. variables [15]. variation [123]. Vereniging [28, 94]. verifiable [177]. verification [131, 131]. versus [98]. Verwendung [259]. Vetter [258]. visual [169]. vitalism [269]. Voor [94, 28]. Voting [287]. vowels [91]. vs [269].

W [138]. Wahrscheinlichkeitsbegriffs [259]. Watanabe [76, 75]. werken [40]. Wetenschappen [28, 94]. Whether [68]. which [194]. whole [191]. Wijsbegeerte [28, 94]. without [20]. Word [224]. work [143]. works [40]. world [100].

zielgerichteter [45]. zum [45]. Zur [58, 259].

References

Anonymous:1960:Ea

[1] Anonymous. Editorial. Synthese, 12(1):5, March 1960. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

     springer.com/article/10.1007/BF00485529.

Rynin:1960:E

[2] David Rynin. Evidence. Synthese, 12(1):6–24, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

     link.springer.com/article/10.1007/BF00485530.

Martin:1960:DSI

[3] Norman M. Martin. Deduction and strict implication. Synthese, 12(1): 25–33, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485531.

Rescher:1960:PTN

[4] Nicholas Rescher. A problem in the theory of numerical estimation. Synthese, 12(1):34–39, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485532.

Dodd:1960:HRI

[5] Stuart C. Dodd. How random interacting organizes a population. Synthese, 12(1):40–70, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485533.

Palmieri:1960:BPN

[6] L. E. Palmieri. Bare particulars, names, and elementary propositions. Synthese, 12(1):71–78, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485534.

Feibleman:1960:PS

[7] James K. Feibleman. The psychology of the scientist. Synthese, 12(1): 79–113, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485535.

Lorenzen:1960:CAM

[8] Paul Lorenzen. Constructive and axiomatic mathematics. Synthese, 12 (1):114–119, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485536.

Anonymous:1960:C

[9] Anonymous. Congresses. Synthese, 12(1):120, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485537.

Anonymous:1960:B

[10] Anonymous. Bookreviews. Synthese, 12(1):120, March 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485538.

Apostel:1960:TFS

[11] Leo Apostel. Towards the formal study of models in the non-formal sciences. Synthese, 12(2–3):125–161, September 1960. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485092.

Atkinson:1960:UME

[12] Richard C. Atkinson. The use of models in experimental psychology. Synthese, 12(2–3):162–171, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485093.

Beth:1960:SPT

[13] E. W. Beth. Semantics of physical theories. Synthese, 12(2–3):172–175, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485094.

Destouches:1960:NMM

[14] Jean-Louis Destouches. Sur la notion de modèle en microphysique. (French) [On the notion of model in microphysics]. Synthese, 12(2–3): 176–181, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485095.

Feys:1960:MVD

[15] Robert Feys. Modèles á variables de différentes sortes pour les logiques modales M′′ ou S5. (French) [Models of variables of different kinds for modal logics M′′ or S5]. Synthese, 12(2–3):182–196, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485096.

Fraisse:1960:MAL

[16] R. Fraïssé. Les modèles et l’algèbre logique. (French) [Models and logical algebra]. Synthese, 12(2–3):197–201, September 1960. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485097.

Freudenthal:1960:MAP

[17] Hans Freudenthal. Models in applied probability. Synthese, 12(2–3):202– 212, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485098.

Frey:1960:SIM

[18] G. Frey. Symbolische und ikonische Modelle. (German) [Symbolic and iconic models]. Synthese, 12(2–3):213–221, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485099.

Groenewold:1960:MP

[19] H. J. Groenewold. The model in physics. Synthese, 12(2–3):222–227, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485100.

Grzegorczyk:1960:AGP

[20] Andrzej Grzegorczyk. Axiomatizability of geometry without points. Synthese, 12(2–3):228–235, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485101.

Guillaume:1960:CAS

[21] Marcel Guillaume. Certains aspects syntactiques d’une notion de modèle: Relativisation d’une fonction logique de choix. (French) [Some syntactic aspects of a model notion: Relativisation of a logical choice function]. Synthese, 12(2–3):236–248, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485102.

Kuipers:1960:MI

[22] A. Kuipers. Model and insight. Synthese, 12(2–3):249–256, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485103.

Mulckhuyse:1960:MM

[23] J. J. Mulckhuyse. Molecules and models. Synthese, 12(2–3):257–275, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485104.

Yem:1960:MIE

[24] Pham Xuân Yêm. Modèle d’interaction entre corpuscules en théorie fonctionnelle. (French) [Model of interaction between corpuscles in functional theory]. Synthese, 12(2–3):276–278, September 1960. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485105.

Staal:1960:FSI

[25] J. F. Staal. Formal structures in Indian logic. Synthese, 12(2–3):279– 286, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485106.

Suppes:1960:CMU

[26] Patrick Suppes. A comparison of the meaning and uses of models in mathematics and the empirical sciences. Synthese, 12(2–3):287–301, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485107.

Ubbink:1960:MDK

[27] J. B. Ubbink. Model, description and knowledge. Synthese, 12(2–3): 302–319, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485108.

Anonymous:1960:NVV

[28] Anonymous. Nederlandse Vereniging voor Logica en Wijsbegeerte der Exacte Wetenschappen. (Dutch) [Dutch Association for Logic and Philosophy of the Exact Sciences]. Synthese, 12(2–3):320–321, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485109.

Anonymous:1960:Eb

[29] Anonymous. Errata. Synthese, 12(2–3):322, September 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485110;   http://link.
springer.com/content/pdf/10.1007/BF00485110.pdf.

Anonymous:1960:EN

[30] Anonymous. Editor’s note. Synthese, 12(4):327, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485419.

Bar-Hillel:1960:PRP

[31] Yehoshua Bar-Hillel. A prerequisite for rational philosophical discussion. Synthese, 12(4):328–332, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485420.

Fraenkel:1960:EL

[32] Abraham A. Fraenkel. Epistemology and logic. Synthese, 12(4):333– 337, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485421.

Jorgensen:1960:SRC

[33] Jørgen Jørgensen. Some remarks concerning languages, calculuses, and logic. Synthese, 12(4):338–349, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485422.

Quine:1960:CLT

[34] W. V. Quine. Carnap and logical truth. Synthese, 12(4):350–374, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485423.

Beth:1960:EI

[35] Evert W. Beth. Extension and intension. Synthese, 12(4):375–379, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485424.

Mehlberg:1960:PSP

[36] Henryk Mehlberg. The present situation in the philosophy of mathematics. Synthese, 12(4):380–414, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485425.

Menger:1960:COR

[37] Karl Menger. A counterpart of Occam’s razor in pure and applied mathematics ontological uses. Synthese, 12(4):415–428, December 1960. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485426.

Montague:1960:TGT

[38] Richard Montague. Towards a general theory of computability. Synthese, 12(4):429–438, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485427.

Hempel:1960:II

[39] Carl G. Hempel. Inductive inconsistencies. Synthese, 12(4):439–469, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485428.

Durr:1960:BAL

[40] Karl Dürr. Beleuchtung von anwendungen der logistik in werken von Rudolf Carnap. (German) [Illumination of applications of logistics in works of Rudolf Carnap]. Synthese, 12(4):470–480, December 1960. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485429.

Naess:1960:TQA

[41] Arne Naess. Typology of questionnaires adopted to the study of expressions with closely related meanings. Synthese, 12(4):481–494, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485430.

Tennessen:1960:PIL

[42] H. Tennessen. Permissible and impermissible locutions. Synthese, 12(4): 495–508, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485431.

Rynin:1960:NCS

[43] David Rynin. Non-cognitive synonymy and the definability of ‘good’. Synthese, 12(4):509–516, December 1960. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485432.

Morris:1960:HIE

[44] Charles Morris. On the history of the international encyclopedia of unified science. Synthese, 12(4):517–521, December 1960. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485433.

Stegmuller:1961:EBP

[45] W. Stegmüller. Einige Beiträge zum Problem der Teleologie und der Analyse von Systemen mit zielgerichteter Organisation. (German) [Some contributions to the problem of teleology and the analysis of systems with a targeted organization]. Synthese, 13(1):5–40, March 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485934.

Goudge:1961:GF

[46] T. A. Goudge. The genetic fallacy. Synthese, 13(1):41–48, March 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485935.

Naess:1961:S

[47] Arne Naess. A study of ‘or’. Synthese, 13(1):49–60, March 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485936.

VanDuijn:1961:MTF

[48] P. Van Duijn. A model for theory finding in science. Synthese, 13(1): 61–67, March 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485937.

Workman:1961:LSP

[49] Rollin W. Workman. The logical status of the principle of induction. Synthese, 13(1):68–74, March 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485938.

Aeschlimann:1961:RCE

[50] Mme Florence Aeschlimann. Le rôle de la convention dans l’espace d’un observateur isolé. (French) [The role of convention in the space of an isolated observer]. Synthese, 13(1):75–85, March 1961. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485939.

Heyting:1961:BR

[51] A. Heyting, Dr. P. H. Esser, and Gabriël Nuchelmans. Book reviews. Synthese, 13(1):86–97, March 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485940.

Smart:1961:GTC

[52] J. J. C. Smart. Gödel’s theorem, Church’s theorem, and mechanism. Synthese, 13(2):105–110, June 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00634578.

Kaiser:1961:CTD

[53] L. Kaiser. The concept of three dimensions on different psychic levels. Synthese, 13(2):111–126, June 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00634579.

Kuhn:1961:TUL

[54] Alfred Kuhn. Toward a uniform language of information and knowledge. Synthese, 13(2):127–153, June 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00634580.

Fellows:1961:SCI

[55] Erwin W. Fellows. Social and cultural influences in the development of science. Synthese, 13(2):154–172, June 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00634581.

Simon:1961:CMU

[56] Herbert A. Simon. Comment: The meaning and uses of models. Synthese, 13(2):173–174, June 1961. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00634582.

Whitmore:1961:BR

[57] P. J. S. Whitmore. Book reviews. Synthese, 13(2):175–178, June 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00634583.

Bernays:1961:RSE

[58] Paul Bernays. Zur Rolle der Sprache in erkenntnistheoretischer Hinsicht. (German) [On the role of language in epistemological theory]. Synthese, 13 (3):185–200, September 1961. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00489883.

Bedau:1961:CQM

[59] Hugo Bedau and Paul Oppenheim. Complementarity in quantum mechanics: A logical analysis. Synthese, 13(3):201–232, September 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00489884.

Tillman:1961:IE

[60] Frank Tillman and B. Roswell Russell. Information and entropy. Synthese, 13(3):233–241, September 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00489885.

Rescher:1961:NDR

[61] Nicholas Rescher. Non-deductive rules of inference and problems in the analysis of inductive reasoning. Synthese, 13(3):242–251, September 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00489886.

Reesinck:1961:BR

[62] J. J. M. Reesinck. Book reviews. Synthese, 13(3):252–253, September 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00489887.

Watanabe:1961:MMB

[63] Satosi Watanabe. A model of mind-body relation in terms of modular logic. Synthese, 13(4):261–302, December 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486628. See comments [75, 76].

Marcus:1961:MIL

[64] Ruth Barcan Marcus. Modalities and intensional languages. Synthese, 13 (4):303–322, December 1961. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486629. See reply [65].

Quine:1961:RPM

[65] W. V. Quine. Reply to Professor Marcus. Synthese, 13(4):323–330, December 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486630. See [64].

Menger:1961:COR

[66] Karl Menger. A counterpart of Occam’s razor in pure and applied mathematics; semantic USES. Synthese, 13(4):331–349, December 1961. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486631.

Frege:1961:ALI

[67] G. Frege. About the law of inertia. Synthese, 13(4):350–363, December 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00486632.

Dutton:1961:W

[68] John D. Dutton. Whether. Synthese, 13(4):364–371, December 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486633.

Bogen:1961:RKH

[69] James Bogen. Remarks on the Kierkegaard–Hegel controversy. Synthese, 13(4):372–389, December 1961. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486634.

Schiller:1962:IQT

[70] Ralph Schiller. Interpretations of the quantum theory. Synthese, 14(1): 5–16, March 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02274801.

Grunbaum:1962:FTT

[71] Adolf Grünbaum. The falsifiability of theories: Total or partial? A contemporary evaluation of the Duhem–Quine thesis. Synthese, 14(1): 17–34, March 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF02274802. See comments [83].

Fenstad:1962:NS

[72] Jens Erik Fenstad. Notes on synonymy. Synthese, 14(1):35–77, March 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02274803.

Mehlberg:1962:CMC

[73] Josephine J. Mehlberg. A classification of mathematical concepts. Synthese, 14(1):78–86, March 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF02274804.

Reesinck:1962:BR

[74] J. J. M. Reesinck. Book reviews. Synthese, 14(1):87, March 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF02274805.

Solomonoff:1962:CDW

[75] R. J. Solomonoff. Comments on Dr. S. Watanabe’s paper. Synthese, 14(2– 3):97–100, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00881986. See [63].

Goodall:1962:CDW

[76] M. Goodall. Comments on Dr. S. Watanabe’s paper. Synthese, 14(2–3): 101–102, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00881987. See [63].

Lenneberg:1962:RLF

[77] Eric H. Lenneberg. The relationship of language to the formation of concepts. Synthese, 14(2–3):103–109, September 1962. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00881988.

Tisza:1962:LSP

[78] Laszlo Tisza. The logical structure of physics. Synthese, 14(2–3):110– 131, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00881989.

Anonymous:1962:DPR

[79] Anonymous. Discussion on the paper of Ruth B. Marcus. Synthese, 14(2– 3):132–143, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00881990.

Anderson:1962:TFA

[80] Alan Ross Anderson and Omar Khayyam Moore. Toward a formal analysis of cultural objects. Synthese, 14(2–3):144–170, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00881991.

Siegel:1962:OAH

[81] Armand Siegel. Operational aspects of hidden-variable quantum theories. Synthese, 14(2–3):171–188, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00881992.

Shimony:1962:CPP

[82] Abner Shimony. Comments on the papers of Prof. S. Schiller and Prof. A. Siegel. Synthese, 14(2–3):189–192, September 1962. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00881993.

Cohen:1962:CGP

[83] Robert S. Cohen. Comments on A. Grünbaum’s paper. Synthese, 14(2–3): 193–195, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00881994. See [71].

Caton:1962:SLT

[84] Charles E. Caton. A stipulation of logical truth in a modal propositional calculus. Synthese, 14(2–3):196–199, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00881995.

Rescher:1962:SRN

[85] Nicholas Rescher. The stochastic revolution and the nature of scientific explanation. Synthese, 14(2–3):200–215, September 1962. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00881996.

vanHolk:1962:T

[86] A. G. F. van Holk. On transformations. Synthese, 14(2–3):216–223, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00881997.

Esser:1962:BR

[87] P. H. Esser and E. W. Beth. Book reviews. Synthese, 14(2–3):224– 228, September 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00881998.

Rand:1962:LDS

[88] Rose Rand. The logic of demand-sentences. Synthese, 14(4):237–254, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869308.

Schock:1962:DUR

[89] Rolf Schock. On determinism, the universe, and related concepts. Synthese, 14(4):255–276, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869309.

Dodd:1962:HML

[90] Stuart C. Dodd. How momental laws can be developed in sociology by deducing testable and predictive “actance” models from transacts. Synthese, 14(4):277–299, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869310.

Kaiser:1962:CAS

[91] L. Kaiser. Communication of affects by single vowels. Synthese, 14(4): 300–319, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869311.

vanThemaat:1962:FAL

[92] W. A. Verloren van Themaat. Formalized and artificial languages. Synthese, 14(4):320–326, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869312.

Reesinck:1962:R

[93] J. J. M. Reesinck. Review. Synthese, 14(4):327, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869313.

Anonymous:1962:NVV

[94] Anonymous. Nederlandse Vereniging Voor Logica En Wijsbegeerte Der Exacte Wetenschappen. (Dutch) [Dutch Association for Logic and Philosophy of Physical Sciences]. Synthese, 14(4):328–329, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869314.

Anonymous:1962:CIU

[95] Anonymous. Communications of the international union of history and philosophy of science division of logic, methodology and philosophy of science. Synthese, 14(4):330–340, December 1962. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869315.

Beth:1963:RBF

[96] Evert W. Beth. The relationship between formalised languages and natural language. Synthese, 15(1):1–16, January 1963. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484839.

Rogers:1963:SFS

[97] Robert Rogers. A survey of formal semantics. Synthese, 15(1):17– 56, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484840.

Kemeny:1963:AVF

[98] John G. Kemeny. Analyticity versus fuzziness. Synthese, 15(1):57– 80, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484841.

Martin:1963:TLI

[99] R. M. Martin. Toward a logic of intentions. Synthese, 15(1):81– 102, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484842.

Quine:1963:STC

[100] Willard van Orman Quine. On simple theories of a complex world. Synthese, 15(1):103–106, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484843.

Rashevsky:1963:DRS

[101] N. Rashevsky. The devious roads of science. Synthese, 15(1):107– 114, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484844.

Blum:1963:CO

[102] Harold F. Blum. Complexity and organization. Synthese, 15(1):115– 121, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484845.

Beth:1963:R

[103] E. W. Beth. Review. Synthese, 15(1):122, January 1963. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484846.

Anonymous:1963:C

[104] Anonymous. Communication. Synthese, 15(1):122, January 1963. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484847.

Lejewski:1963:ASE

[105] CzesLaw Lejewski. Aristotle’s syllogistic and its extensions. Synthese, 15 (1):125–154, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484848.

Fitch:1963:ASR

[106] Frederic B. Fitch. Algebraic simplification of redundant sequential circuits. Synthese, 15(1):155–166, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484849.

Popper:1963:CNC

[107] K. R. Popper. Creative and non-creative definitions in the calculus of probability. Synthese, 15(1):167–186, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484850.

Alexander:1963:ST

[108] Peter Alexander. Speculations and theories. Synthese, 15(1):187–203, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484851.

Mays:1963:PMT

[109] W. Mays. Probability models and thought and learning processes. Synthese, 15(1):204–221, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484852.

Lewontin:1963:MMM

[110] R. C. Lewontin. Models, mathematics and metaphors. Synthese, 15(1): 222–244, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484853.

Helmer:1963:GTA

[111] Olaf Helmer. The game-theoretical approach to organization theory. Synthese, 15(1):245–253, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484854.

Dunham:1963:DNS

[112] B. Dunham, D. Fridshal, R. Fridshal, and J. H. North. Design by natural selection. Synthese, 15(1):254–259, January 1963. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484855.

Beckner:1963:MPD

[113] Morton Beckner. Metaphysical presuppositions and the description of biological systems. Synthese, 15(1):260–274, January 1963. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484856.

Bonner:1963:AB

[114] John Tyler Bonner. Analogies in biology. Synthese, 15(1):275–279, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484857.

Martin:1963:MLC

[115] R. M. Martin. On maximum logical candor and extensionality. Synthese, 15(1):283–291, January 1963. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484858.

Luchins:1963:TPS

[116] Abraham S. Luchins and Edith H. Luchins. Two philosophies of science: A study in contrasts. Synthese, 15(1):292–316, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484859.

Caws:1963:QTC

[117] Peter Caws. A quantum theory of causality. Synthese, 15(1):317– 326, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484860.

Peters:1963:RCT

[118] Franz Peters. Russell on class theory. Synthese, 15(1):327–335, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484861.

Muller-Markus:1963:PAR

[119] Siegfried Müller-Markus. Die Prinzipien Der Allgemeinen Relativitätstheorie. (German) [The principles of General Relativity Theory]. Synthese, 15(1):336–374, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484862.

Crow:1963:SRI

[120] C. Crow. Some remarks on induction. Synthese, 15(1):379–388, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484863.

deMott:1963:SRR

[121] Benjamin de Mott. Science and the rejection of realism in art. Synthese, 15(1):389–400, January 1963. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484864.

Margolis:1963:POM

[122] Joseph Margolis. The problem of other minds. Synthese, 15(1):401– 411, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484865.

Sacksteder:1963:SVS

[123] William Sacksteder. Structural variation in science. Synthese, 15(1): 412–423, January 1963. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484866.

Feibleman:1963:MOR

[124] James K. Feibleman. Material objects and the reference of signs. Synthese, 15(1):424–435, January 1963. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484867.

Ackermann:1963:ISS

[125] Robert Ackermann. Inductive simplicity in special cases. Synthese, 15 (1):436–444, January 1963. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484868.

Anonymous:1966:E

[126] Anonymous. Editorial. Synthese, 16(1):1–3, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873570.

vonWright:1966:MA

[127] G. H. von Wright. Memorial address. Synthese, 16(1):4, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873571.

Destouches:1966:A

[128] J. L. Destouches. Allocution. Synthese, 16(1):5–6, September 1966. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873572.

Nieland:1966:BTM

[129] J. J. F. Nieland. Beth’s tableau-method. Synthese, 16(1):7–26, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00873573.

Guillaume:1966:QRT

[130] Marcel Guillaume. Quelques remarques sur les ‘Tableaux de Beth’. (French) [A few remarks on ’Beth’s Tables’]. Synthese, 16(1):27–33, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873574.

Fraisse:1966:HER

[131] Roland Fraïssé. Une hypothèse sur l’extension des relations finies et sa vérification dans certaines classes particulières. (French) [A hypothesis on the extension of finite relations and its verification in certain particular classes]. Synthese, 16(1):34–46, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873575.

Heyting:1966:RTI

[132] A. Heyting. Remarques sur la théorie intuitionniste des espaces linéaires. (French) [Remarks on the intuitionistic theory of linear spaces]. Synthese, 16(1):47–52, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873576.

Mooij:1966:PGH

[133] J. J. A. Mooij. La philosophie géométrique de Henri Poincaré. (French) [The geometric philosophy of Henri Poincaré]. Synthese, 16(1):53–65, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873577.

Destouches:1966:LTP

[134] J. L. Destouches. Logique et théorie physique. (French) [Logic and physical theory]. Synthese, 16(1):66–73, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873578.

Suppes:1966:LPP

[135] Patrick Suppes. L’argument probabiliste pour une logique non classique de la mécanique quantique. (French) [The probabilistic argument for an unconventional logic of quantum mechanics]. Synthese, 16(1):74– 85, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873579.

Heyting:1966:C

[136] A. Heyting. Conclusions. Synthese, 16(1):86–88, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873580.

Anonymous:1966:EN

[137] Anonymous. Editorial note. Synthese, 16(1):89, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873581.

Staal:1966:BWB

[138] J. F. Staal. Bibliography of E. W. Beth. Synthese, 16(1):90–106, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873582.

Juhos:1966:R

[139] Béla Juhos and Howard Smokler. Reviews. Synthese, 16(1):107–122, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873583.

Anonymous:1966:BRa

[140] Anonymous. Books received. Synthese, 16(1):123–124, September 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873584.

Adams:1966:NPM

[141] Ernest W. Adams. On the nature and purpose of measurement. Synthese, 16(2):125–169, November 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485355.

Rozeboom:1966:STN

[142] William W. Rozeboom. Scaling theory and the nature of measurement. Synthese, 16(2):170–233, November 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485356.

Levi:1966:RWP

[143] Isaac Levi. Recent work in probability and induction (review article). Synthese, 16(2):234–244, November 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485357.

Anonymous:1966:CIU

[144] Anonymous. Communications of the international union of history and philosophy of science division of logic, methodology and philosophy of science. Synthese, 16(2):245–252, November 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485358.

Arrow:1966:ETC

[145] Kenneth J. Arrow. Exposition of the theory of choice under uncertainty. Synthese, 16(3–4):253–269, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485082.

Sneed:1966:SLD

[146] Joseph D. Sneed. Strategy and the logic of decision. Synthese, 16(3–4): 270–283, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485083.

Suppes:1966:SFM

[147] Patrick Suppes. Some formal models of grading principles. Synthese, 16(3– 4):284–306, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485084.

Hallden:1966:PLT

[148] Sören Halldén. Preference logic and theory choice. Synthese, 16(3–4): 307–320, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485085.

Chisholm:1966:IPP

[149] Roderick M. Chisholm and Ernest Sosa. Intrinsic preferability and the problem of supererogation. Synthese, 16(3–4):321–331, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485086.

Rescher:1966:NPU

[150] Nicholas Rescher. Notes on preference, utility, and cost. Synthese, 16(3– 4):332–343, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485087.

Hanna:1966:NAF

[151] Joseph F. Hanna. A new approach to the formulation and testing of learning models. Synthese, 16(3–4):344–380, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485088.

OHair:1966:TIT

[152] S. G. O’Hair. Two introductions to the theory of knowledge (review article). Synthese, 16(3–4):381–393, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485089.

Runciman:1966:R

[153] W. G. Runciman and Gabriël Nuchelmans. Reviews. Synthese, 16(3–4): 394–399, December 1966. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485090.

Anonymous:1966:BRb

[154] Anonymous. Books received. Synthese, 16(3–4):400, December 1966. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485091.

Anonymous:1967:EN

[155] Anonymous. Editorial note. Synthese, 17(1):1, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485012.

Chomsky:1967:RCT

[156] Noam Chomsky. Recent contributions to the theory of innate ideas. Synthese, 17(1):2–11, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485013.

Putnam:1967:IHE

[157] Hilary Putnam. The ‘innateness hypothesis’ and explanatory models in linguistics. Synthese, 17(1):12–22, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485014.

Goodman:1967:EA

[158] Nelson Goodman. The epistemological argument. Synthese, 17(1):23– 28, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485015.

Sachs:1967:EMG

[159] Mendel Sachs. On the elementarity of measurement in general relativity: Toward a general theory. Synthese, 17(1):29–53, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485016.

Buchdahl:1967:SSC

[160] Gerd Buchdahl. Semantic sources of the concept of law. Synthese, 17 (1):54–74, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485017.

Toulmin:1967:CRS

[161] Stephen Toulmin. Conceptual revolutions in science. Synthese, 17(1): 75–91, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00485018. See comment [162].

Mink:1967:CST

[162] Louis O. Mink. Comment on Stephen Toulmin’s ‘Conceptual revolutions in science’. Synthese, 17(1):92–99, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485019. See [161].

Roberts:1967:SPA

[163] George W. Roberts. Some points about discontinuity. Synthese, 17(1): 100–103, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485020.

Harre:1967:MPD

[164] R. Harré. More points on discontinuity. Synthese, 17(1):104–106, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485021.

Gellner:1967:R

[165] Ernest Gellner and Hermann Kramer. Reviews. Synthese, 17(1):107– 116, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485022.

Ayer:1967:ARS

[166] A. J. Ayer. Has Austin refuted the sense-datum theory? Synthese, 17 (1):117–140, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485023.

Ellis:1967:PM

[167] Brian Ellis. Physical monism. Synthese, 17(1):141–161, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485024.

Gibson:1967:NRR

[168] James J. Gibson. New reasons for realism. Synthese, 17(1):162–172, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485025.

Roberts:1967:SPG

[169] Fred S. Roberts and Patrick Suppes. Some problems in the geometry of visual perception. Synthese, 17(1):173–201, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485026.

Swartz:1967:CCD

[170] Robert J. Swartz. Color concepts and dispositions. Synthese, 17(1):202– 222, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485027.

Yolton:1967:R

[171] John W. Yolton and K. von Fieandt. Reviews. Synthese, 17(1):223– 232, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485028.

Wittgenstein:1967:BF

[172] Ludwig Wittgenstein. Bemerkungen Über Frazers. (German) [Remark on Frazer]. Synthese, 17(1):233–253, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485029.

Stenius:1967:MLG

[173] Erik Stenius. Mood and language-game. Synthese, 17(1):254–274, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485030. See

comments [174].

Follesdal:1967:CSM

[174] Dagfinn Føllesdal. Comments on Stenius’s ‘Mood and language-game’. Synthese, 17(1):275–280, January 1967. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485031. See [173].

Aaqvist:1967:SPC

[175] Lennart Åqvist. Semantic and pragmatic characterizability of linguistic usage. Synthese, 17(1):281–291, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485032.

Margolis:1967:EA

[176] Joseph Margolis. Entitled to assert. Synthese, 17(1):292–298, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485033.

OHair:1967:PSV

[177] S. G. O’Hair. Performatives and sentences verifiable by their use. Synthese, 17(1):299–303, January 1967. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485034.

Davidson:1967:TM

[178] Donald Davidson. Truth and meaning. Synthese, 17(1):304–323, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485035.

VanHeijenoort:1967:LCL

[179] Jean Van Heijenoort. Logic as calculus and logic as language. Synthese, 17(1):324–330, January 1967. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485036.

Cohen:1967:MAS

[180] A. Cohen. On methods in the analysis of speech perception. Synthese, 17(1):331–343, January 1967. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485037.

Mates:1967:R

[181] Benson Mates, Ryszard Wójcicki, Bruce Vermazen, and Rollin W. Workman. Reviews. Synthese, 17(1):344–367, January 1967. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485038.

Anonymous:1967:BRa

[182] Anonymous. Books received. Synthese, 17(1):368, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485039.

Levi:1967:II

[183] Isaac Levi. Information and inference. Synthese, 17(1):369–391, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485040.

Sneed:1967:EID

[184] Joseph D. Sneed. Entropy, informantion, and decision. Synthese, 17(1): 392–407, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485041.

Szaniawski:1967:VPI

[185] Klemens Szaniawski. The value of perfect information. Synthese, 17(1): 408–424, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485042.

Bergstrom:1967:NM

[186] R. M. Bergström. Neural macrostates. Synthese, 17(1):425–443, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485043.

Caws:1967:R

[187] Peter Caws. Reviews. Synthese, 17(1):444–462, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485044.

Anonymous:1967:BRb

[188] Anonymous. Books received. Synthese, 17(1):463–464, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485045.

Anonymous:1967:C

[189] Anonymous. Correction. Synthese, 17(1):464, January 1967. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485046;   https://link.
springer.com/content/pdf/10.1007/BF00485046.pdf.

Fenstad:1968:SLP

[190] Jens Erik Fenstad. The structure of logical probabilities. Synthese, 18(1): 1–23, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869595.

Jeffrey:1968:WT

[191] Richard C. Jeffrey. The whole truth. Synthese, 18(1):24–27, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869596.

Tornebohm:1968:CHA

[192] Håkan Törnebohm. On the confirmation of hypotheses about regions of existence. Synthese, 18(1):28–45, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869597.

Ruddick:1968:CC

[193] William Ruddick. Causal connection. Synthese, 18(1):46–67, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869598.

Hillinger:1968:GPC

[194] Claude Hillinger. A generalization of the principle of causality, which makes it applicable to evolutionary systems. Synthese, 18(1):68–74, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869599.

Havas:1968:CRT

[195] Peter Havas. Causality requirements and the theory of relativity. Synthese, 18(1):75–102, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869600.

Agassi:1968:NMD

[196] Joseph Agassi. No more discovery in physics? Synthese, 18(1):103– 108, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869601.

Essler:1968:R

[197] W. K. Essler and Abner Shimony. Reviews. Synthese, 18(1):109– 120, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869602.

Anonymous:1968:CIU

[198] Anonymous. Communications of the international union of history and philosophy of science division of logic, methodology and philosophy of science. Synthese, 18(1):121–129, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869603.

Anonymous:1968:BRa

[199] Anonymous. Books received. Synthese, 18(1):130, January 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869604.

Hintikka:1968:E

[200] J. Hintikka. Editorial. Synthese, 18(2–3):131, April 1968. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413772.

Cohen:1968:EMP

[201] Robert S. Cohen. Ernst Mach: Physics, perception and the philosophy of science. Synthese, 18(2–3):132–170, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413773.

Capek:1968:EMB

[202] Milic Capek. Ernst Mach’s biological theory of knowledge. Synthese, 18 (2–3):171–191, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413774.

Brush:1968:MA

[203] S. G. Brush. Mach and atomism. Synthese, 18(2–3):192–215, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413775.

Koslow:1968:MCM

[204] A. Koslow. Mach’s concept of mass: Program and definition. Synthese, 18(2–3):216–233, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413776.

Yourgrau:1968:DEM

[205] Wolfgang Yourgrau and Alwyn van der Merwe. Did Ernst Mach ‘miss the target’? Synthese, 18(2–3):234–250, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413777.

Strauss:1968:ETC

[206] M. Strauss. Einstein’s theories and the critics of Newton. Synthese, 18 (2–3):251–284, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413778.

Thiele:1968:BDP

[207] Joachim Thiele. Briefe Deutscher Philosophen An Ernst Mach. (German) [Letters of a German philosopher to Ernst Mach]. Synthese, 18(2–3): 285–301, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413779.

Mostowski:1968:R

[208] Andrzej Mostowski. Review. Synthese, 18(2–3):302–305, April 1968. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413780.

Anonymous:1968:Aa

[209] Anonymous. Abstracts. Synthese, 18(2–3):306–309, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413781.

Anonymous:1968:R

[210] Anonymous. Rectification. Synthese, 18(2–3):309, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413782;   http://link.
springer.com/content/pdf/10.1007/BF00413782.pdf.

Anonymous:1968:BRb

[211] Anonymous. Books received. Synthese, 18(2–3):310, April 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413783.

Roberts:1968:ATE

[212] Fred S. Roberts and R. Duncan Luce. Axiomatic thermodynamics and extensive measurement. Synthese, 18(4):311–326, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484975.

Duistermaat:1968:EER

[213] J. J. Duistermaat. Energy and entropy as real morphisms for addition and order. Synthese, 18(4):327–393, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484976.

Wilson:1968:OUT

[214] Fred Wilson. Is operationism unjust to temperature? Synthese, 18(4): 394–422, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484977.

Hansson:1968:FAP

[215] Bengt Hansson. Fundamental axioms for preference relations. Synthese, 18(4):423–442, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484978.

Hansson:1968:CSP

[216] Bengt Hansson. Choice structures and preference relations. Synthese, 18 (4):443–458, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484979.

Kordig:1968:PD

[217] Carl R. Kordig. On prescribing description. Synthese, 18(4):459–461, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484980.

Toulmin:1968:R

[218] Stephen Toulmin. Reply. Synthese, 18(4):462–463, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484981.

Sneed:1968:R

[219] Joseph D. Sneed. Review. Synthese, 18(4):464–467, October 1968. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484982.

Anonymous:1968:Ab

[220] Anonymous. Abstracts. Synthese, 18(4):468–473, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484983.

Anonymous:1968:BRc

[221] Anonymous. Books received. Synthese, 18(4):474, October 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484984.

Davidson:1968:EI

[222] Donald Davidson and Jaakko Hintikka. Editorial introduction. Synthese, 19(1–2):1–2, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00568045.

Smart:1968:QPS

[223] J. J. C. Smart. Quine’s philosophy of science. Synthese, 19(1–2):3– 13, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568046.

Harman:1968:ITM

[224] Gilbert Harman. An introduction to ‘translation and meaning’ chapter two of word and object. Synthese, 19(1–2):14–26, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00568047.

Stenius:1968:BOT

[225] Erik Stenius. Beginning with ordinary things. Synthese, 19(1–2):27– 52, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568048.

Chomsky:1968:QEA

[226] Noam Chomsky. Quine’s empirical assumptions. Synthese, 19(1–2):53– 68, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568049.

Hintikka:1968:BCR

[227] Jaakko Hintikka. Behavioral criteria of radical translation. Synthese, 19 (1–2):69–81, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00568050.

Stroud:1968:CIT

[228] Barry Stroud. Conventionalism and the indeterminacy of translation. Synthese, 19(1–2):82–96, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00568051.

Strawson:1968:STP

[229] P. F. Strawson. Singular terms and predication. Synthese, 19(1–2):97– 117, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568052.

Geach:1968:QSI

[230] P. T. Geach. Quine’s syntactical insights. Synthese, 19(1–2):118–129, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568053.

Davidson:1968:S

[231] Donald Davidson. On saying that. Synthese, 19(1–2):130–146, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00568054.

Follesdal:1968:QM

[232] Dagfinn Føllesdal. Quine on modality. Synthese, 19(1–2):147–157, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568055.

Sellars:1968:SPA

[233] Wilfrid Sellars. Some problems about belief. Synthese, 19(1–2):158– 177, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568056.

Kaplan:1968:Q

[234] David Kaplan. Quantifying in. Synthese, 19(1–2):178–214, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00568057.

Berry:1968:LP

[235] George Berry. Logic with Platonism. Synthese, 19(1–2):215–249, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00568058.

Jensen:1968:CSM

[236] Ronald Björn Jensen. On the consistency of a slight (?) modification of Quine’s mew foundations. Synthese, 19(1–2):250–263, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00568059.

Quine:1968:R

[237] W. V. Quine. Replies. Synthese, 19(1–2):264–322, December 1968. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00568060.

Anonymous:1968:BRd

[238] Anonymous. Books received. Synthese, 19(1–2):323–324, December 1968. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00568061.

Goguen:1969:LIC

[239] J. A. Goguen. The logic of inexact concepts. Synthese, 19(3–4):325– 373, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485654.

Przelecki:1969:PA

[240] Marian Przelecki and Ryszard Wójcicki. The problem of analyticity. Synthese, 19(3–4):374–399, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485655.

Yolton:1969:GR

[241] John W. Yolton. Gibson’s realism. Synthese, 19(3–4):400–407, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485656.

Gibson:1969:TSQ

[242] James J. Gibson. Are there sensory qualities of objects? Synthese, 19 (3–4):408–409, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485657.

Arbini:1969:CLC

[243] Ronald Arbini. Comments on linguistic competence and language acquisition. Synthese, 19(3–4):410–424, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485658.

Harman:1969:RA

[244] Gilbert Harman. Reply to arbini. Synthese, 19(3–4):425–432, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485659.

Strauss:1969:CBF

[245] M. Strauss. Corrections to Bunge’s foundations of physics (1967). Synthese, 19(3–4):433–442, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485660;   http://link.springer.com/content/
pdf/10.1007/BF00485660.pdf.

Bunge:1969:CFP

[246] Mario Bunge. Corrections to foundations of physics: Correct and incorrect. Synthese, 19(3–4):443–452, April 1969. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485661;   http://link.
springer.com/content/pdf/10.1007/BF00485661.pdf.

Agassi:1969:COB

[247] Joseph Agassi. Changing our background-knowledge. Synthese, 19(3–4): 453–464, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485662.

Anonymous:1969:ENa

[248] Anonymous. Editorial note. Synthese, 19(3–4):465, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485663.

Hintikka:1969:R

[249] Jaakko Hintikka, Margaret Donaldson, and Robert J. Fogelin. Reviews. Synthese, 19(3–4):466–480, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485664.

Anonymous:1969:APSa

[250] Anonymous. Abstracts from Philosophy of Science. Synthese, 19(3–4): 481–485, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485665.

Anonymous:1969:BRa

[251] Anonymous. Books received. Synthese, 19(3–4):486–487, April 1969. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485666.

Anonymous:1969:ENb

[252] Anonymous. Editorial note. Synthese, 20(1):1, June 1969. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567231.

DeFinetti:1969:IPP

[253] Bruno De Finetti. Initial probabilities: A prerequisite for any valid induction. Synthese, 20(1):2–16, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567232.

Good:1969:DBF

[254] I. J. Good. Discussion of Bruno de Finetti’s paper ‘Initial probabilities: A prerequisite for any valid induction’. (French) [Discussion of Bruno of Finetti’s paper ‘Initial probabilities: A prerequisite for any valid induction’]. Synthese, 20(1):17–24, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567233.

Hacking:1969:LII

[255] Ian Hacking. Linguistically invariant inductive logic. Synthese, 20(1): 25–47, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567234.

Levi:1969:CLI

[256] Isaac Levi. Confirmation, linguistic invariance, and conceptual innovation. Synthese, 20(1):48–55, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567235.

Vetter:1969:LPM

[257] Hermann Vetter. Logical probability, mathematical statistics, and the problem of induction. Synthese, 20(1):56–71, June 1969. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567236.

Hintikka:1969:SIL

[258] Jaakko Hintikka. Statistics, induction, and lawlikeness: Comments on Dr. Vetter’s paper. Synthese, 20(1):72–83, June 1969. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567237.

VonKutschera:1969:PNV

[259] Franz Von Kutschera. Zur Problematik der naturwissenschaftlichen Verwendung des subjektiven Wahrscheinlichkeitsbegriffs. (German) [On the problem of the scientific use of the subjective probability concept]. Synthese, 20(1):84–103, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567238.

Frey:1969:GGS

[260] Gerhard Frey. Über die Gültigkeit genereller Sätze. (German) [On the validity of general sentences]. Synthese, 20(1):104–120, June 1969. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567239.

Baumer:1969:DPT

[261] William H. Baumer. In defense of a principal theorem. Synthese, 20(1): 121–142, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567240.

Levi:1969:R

[262] Isaac Levi and B. Juhos. Reviews. Synthese, 20(1):143–153, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567241.

Anonymous:1969:ABJ

[263] Anonymous. Abstracts from the British Journal for the Philosophy of Science. Synthese, 20(1):154–156, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567242.

Hull:1969:WPB

[264] David L. Hull. What philosophy of biology is not. Synthese, 20(2):157– 184, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413784.

Manier:1969:EMB

[265] Edward Manier. The experimental method in biology. Synthese, 20(2): 185–205, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413785.

Manier:1969:FSE

[266] Edward Manier. ‘Fitness’ and some explanatory patterns in biology. Synthese, 20(2):206–218, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413786.

Ackermann:1969:MMB

[267] Robert Ackermann. Mechanism, methodology, and biological theory. Synthese, 20(2):219–229, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413787.

Causey:1969:PSR

[268] Robert L. Causey. Polanyi on structure and reduction. Synthese, 20 (2):230–237, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413788.

Hein:1969:MBV

[269] Hilde Hein. Molecular biology vs. organicism: The enduring dispute between mechanism and vitalism. Synthese, 20(2):238–253, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413789.

Lagerspetz:1969:ICB

[270] Kari Y. H. Lagerspetz. Individuality and creativity: Is biology different? Synthese, 20(2):254–260, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413790.

Macklin:1969:TBS

[271] Martin Macklin and Ruth Macklin. Theoretical biology: A statement and defense. Synthese, 20(2):261–276, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413791.

Roll-Hansen:1969:RBP

[272] Nils Roll-Hansen. On the reduction of biology to physical science. Synthese, 20(2):277–289, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413792.

Anonymous:1969:APSb

[273] Anonymous. Abstracts from Philosophy of Science. Synthese, 20(2):290– 292, August 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413793.

Anonymous:1969:BRb

[274] Anonymous. Books received. Synthese, 20(2):293–294, August 1969. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413794.

Anonymous:1969:ENc

[275] Anonymous. Editorial note. Synthese, 20(3):295, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413730.

Levi:1969:SHC

[276] Isaac Levi. Are statistical hypotheses covering laws? Synthese, 20(3): 297–307, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413731.

Hanna:1969:EPD

[277] Joseph F. Hanna. Explanation, prediction, description, and information theory. Synthese, 20(3):308–334, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413732.

Rosenkrantz:1969:E

[278] Roger D. Rosenkrantz. On explanation. Synthese, 20(3):335–370, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413733.

Giere:1969:BSB

[279] Ronald N. Giere. Bayesian statistics and biased procedures. Synthese, 20 (3):371–387, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413734.

Macklin:1969:EAR

[280] Ruth Macklin. Explanation and action: Recent issues and controversies. Synthese, 20(3):388–415, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413735.

Stouland:1969:R

[281] Frederick Stouland. Review. Synthese, 20(3):416–425, October 1969. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413736.

Anonymous:1969:BRc

[282] Anonymous. Books received. Synthese, 20(3):426, October 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413737.

Wold:1969:MEP

[283] Herman O. Wold. Mergers of economics and philosophy of science. Synthese, 20(4):427–482, December 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869929.

Marschak:1969:ET

[284] Jacob Marschak. On econometric tools. Synthese, 20(4):483–488, December 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869930.

Fisher:1969:CSE

[285] Franklin M. Fisher. Causation and specification in economic theory and econometrics. Synthese, 20(4):489–500, December 1969. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869931.

Brunner:1969:ACQ

[286] Karl Brunner. ‘Assumptions’ and the cognitive quality of theories. Synthese, 20(4):501–525, December 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869932.

Hansson:1969:VGD

[287] Bengt Hansson. Voting and group decision functions. Synthese, 20(4): 526–537, December 1969. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869933.

Anonymous:1969:BRd

[288] Anonymous. Books received. Synthese, 20(4):538, December 1969. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869934.